OpenStudy (anonymous):

simplify f(x) = 3x^2-4x+1/3x-1 Using complete sentences, explain why f(1) = 0, f(0) = -1 and f(-1) = -2 yet f( 1/3 ) is undefined. Make sure to show your work.

OpenStudy (whpalmer4):

\[f(x) = \frac{3x^2-4x+1 } { 3x-1}\] I would start by factoring the numerator. To answer the second part, what happens when \(x = \frac{1}{3}\)?

OpenStudy (anonymous):

i am not sure

OpenStudy (whpalmer4):

Okay, let's look at the denominator: \(3x-1\) What do you get if you plug in \(x = \frac{1}{3}\) in that?